### v1

# VidLPRO: A Video-Language Pre-training Framework for Robotic and Laparoscopic Surgery

Abstract
We introduce VidLPRO, a novel video-language (VL) pre-training framework designed specifically for robotic and laparoscopic surgery. While existing surgical VL models primarily rely on contrastive learning, we propose a more comprehensive approach to capture the intricate temporal dynamics and align video with language. VidLPRO integrates video-text contrastive learning, video-text matching, and masked language modeling objectives to learn rich VL representations. To support this framework, we present GenSurg+, a carefully curated dataset derived from GenSurgery, comprising 17k surgical video clips paired with captions generated by GPT-4 using transcripts extracted by the Whisper model. This dataset addresses the need for large-scale, high-quality VL data in the surgical domain. Extensive experiments on benchmark datasets, including Cholec80 and AutoLaparo, demonstrate the efficacy of our approach. VidLPRO achieves state-of-the-art performance in zero-shot surgical phase recognition, significantly outperforming existing surgical VL models such as SurgVLP and HecVL. Our model demonstrates improvements of up to 21.5% in accuracy and 15.7% in F1 score, setting a new benchmark in the field. Notably, VidLPRO exhibits robust performance even with single-frame inference, while effectively scaling with increased temporal context. Ablation studies reveal the impact of frame sampling strategies on model performance and computational efficiency. These results underscore VidLPRO’s potential as a foundation model for surgical video understanding.
1 Introduction
The field of surgical computer vision has seen significant advancements in recent years, driven by the growing demand for artificial intelligence (AI) applications in healthcare. A notable increase in research has led to the development of deep learning models that enable surgical workflow recognition [1, 2, 3], enhance surgical scene understanding [4, 5, 6] and reconstruction [7, 8, 9]. As surgical procedures grow more complex and technology-driven, the demand for intelligent systems that support surgeons throughout the entire surgical journey - from preoperative planning to intra-operative guidance and post-operative analysis [10] - becomes increasingly crucial for enhancing patient outcomes, streamlining workflows, and enhance overall surgical efficiency [11, 12].
Despite these promising applications, the development and implementation of these systems in surgical domain face several challenges. One of the primary challenges is the complexity and variability inherent in surgical procedures. Unlike many standardized video datasets, surgical videos capture highly dynamic environments where the visual content can vary significantly based on the specific procedure, patient anatomy, surgeon technique, and unexpected complications [13, 14]. This variability makes it difficult to develop robust models that can generalize across different surgical scenarios. Another significant challenge is the scarcity of large-scale annotated surgical datasets. Unlike in other domains where data can be more readily collected and labeled, surgical data is subject to strict privacy regulations and requires expert annotation, which is both time-consuming and expensive [11, 15]. This limitation hinders the development of data-hungry deep learning models and necessitates innovative approaches to leverage limited labeled data effectively. The long duration of surgical procedures also poses a unique challenge. Surgical videos often span several hours, requiring models to capture and process long-range temporal dependencies [10]. This is in stark contrast to many general video understanding tasks that typically deal with short clips lasting only a few seconds or minutes. Furthermore, interpreting surgical videos requires specialized medical knowledge, making it challenging to apply general-purpose video understanding models directly to surgical tasks [12].Lastly, the fine-grained nature of surgical actions and the subtle visual cues that distinguish different phases or steps of a procedure add another layer of complexity. Models must be capable of detecting and interpreting small but crucial details in the surgical field, often in the presence of occlusions, reflections, and rapid camera movements [16, 17].
Recently, Multimodal learning, which integrates multiple modalities such visual data, text data, audio, depth maps etc., has emerged as a viable strategy in computer vision domain. Specifically, Vision-Language Pre-training (VLP) which leverages large-scale datasets of paired visual and free-from textual data, can reduce the reliance on annotated datasets, enabling more efficient and effective learning. It enables models to learn rich and generalizable representations that can be adapted to various downstream tasks with minimal fine-tuning such as image-text retrieval [18, 19, 20], visual question answering [21, 22, 23, 24, 25], video understanding [26, 27, 28, 29, 30] and zero-shot classification [31, 32]. The potential of VLP to capture complex relationships between visual content and natural language descriptions makes it particularly appealing for the surgical domain, where procedures are often accompanied by detailed textual reports or narrations.
Recent efforts have begun to explore the application of VLP techniques to surgical video analysis. Notable approaches include SurgVLP [33], which leverages surgical video lectures and their transcripts to learn multi-modal representations, and HecVL [34], which proposes a hierarchical pre-training framework for zero-shot surgical phase recognition. While these methods have shown promising results, they still face several limitations. A significant challenge has been the lack of large-scale, diverse datasets for surgical VLP. The introduction of the GenSurgery dataset [35] was a step forward, providing a substantial collection of surgical videos. However, this dataset had limitations, including a lack of paired textual data, inconsistent audio quality, and the presence of non-informative content. Our GenSurg+ dataset addresses these issues by rigorously filtering the original data, adding high-quality captions, and ensuring rich linguistic context. Despite this progress, existing approaches still struggle with insufficient temporal modeling, failing to capture long-range dependencies in surgical videos effectively. Many current methods show reduced performance when applied to new surgical procedures or tasks not seen during pre-training, indicating limited generalization capabilities. Additionally, most approaches rely solely on video-text contrastive (VTC) learning as shown in Figure 1, missing out on the benefits of other pretraining objectives that could enhance the model’s understanding of surgical content and context. Addressing these limitations is crucial for advancing the field of surgical VLP and developing more robust and versatile models for surgical video understanding.
Figure 1: Current approaches (left) rely on video-text contrastive loss only, while our method (right), besides contrastive loss, employ video-text matching loss and masked language modeling to enhance cross-modal fusion and surgical language.
To address the limitations of existing surgical VLP approaches, we present VidLPRO and GenSurg+, a novel framework and dataset for robotic and laparoscopic surgical video-language foundation models. VidLPRO builds upon recent advancements in video-language pre-training, incorporating a Vision Transformer (ViT) as the video encoder, BERT as the text encoder, and a multimodal fusion module. Our model employs a combination of Video-Text Contrastive Learning (VTC), Video-Text Matching (VTM), and Masked Language Modeling (MLM) objectives to learn nuanced, context-aware representations of surgical procedures as shown in Figure 1. We also introduce GenSurg+, an enhanced version of the GenSurgery dataset [35], containing 17k 45-second clips of endoscopic robotic surgery with high quality captions generated using raw narration and GPT-4. In zero-shot surgical phase recognition, VidLPRO significantly outperforms the current state-of-the-art on both Cholec80 [14] and AutoLaparo [36]. More specifically, on Cholec80, it achieves 57.1% accuracy and 32.1% F1 score, surpassing HecVL by 15.4% and 5.8% respectively. Our ablation studies demonstrate VidLPRO’s robustness across different frame sampling rates, with performance scaling effectively as frame count increases. These results highlight the effectiveness of our pre-training approach, the quality of GenSurg+, and VidLPRO’s potential to generalize across different surgical procedures and tasks, crucial for developing adaptive AI systems for diverse surgical environments.
2 Related Work
2.1 Vision-Language models
Most of the vision-language approaches can be categorized into two groups. The first group focuses on training multi-modal encoders [37, 38, 39, 40, 41] while second group focuses on training uni-modal vision and text encoders [31, 32, 42, 43]. In context of surgical domain, Surgical-VQA [44], SurgicalGPT [45] propose vision-language model for visual question answering. Surgical-LVLM [46] adapts large vision-language model by introducing specialized Visual Perception LoRA blocks for grounded visual question answering in robotic surgery.
2.2 Surgical Video-Language Pretraining
The application of video-language pre-training (VLP) techniques to the surgical domain is a recent development that shows great promise for advancing surgical video analysis. Two notable approaches in this emerging field are SurgVLP [33] and HecVL [34], which have made significant strides in adapting VLP methods to the unique challenges of surgical data. SurgVLP [33] uses contrastive learning objective to learn multi-modal representations from surgical video lectures. This method leverages a large dataset of surgical videos paired with transcribed audio, using multiple complementary automatic speech recognition (ASR) systems to generate text annotations. Building upon this foundation, HecVL [34] proposes a hierarchical video-language pre-training framework specifically designed for zero-shot surgical phase recognition. This approach addresses the challenge of capturing both fine-grained actions and high-level surgical concepts by incorporating hierarchical textual supervision. VidLPRO, on the other-hand, introduces multiple pre-training objective beyond mere contrastive learning to capture more rich multi-modal representations.
2.3 Surgical Phase Recognition
Surgical phase recognition aims to automatically identify and segment different stages of a surgical procedure. Traditional approaches to surgical phase recognition often relied on hand-crafted features and classical machine learning techniques [47, 2]. However, with the advent of deep learning, there has been a shift towards more sophisticated models that can automatically learn relevant features from raw video data [13, 48]. Following these, many one-stage approaches [49, 50, 51] have been proposed to learn spatio-temporal features. However, one-stage approaches fail to capture the long-term spatial-temporal dependency. To address this limitation, two-stage solutions [52, 53, 54, 55] are proposed which first extract the spatial or spatio-temporal features using the feature extractor and then employ a temporal model on the top of these features to learn long-term dependency. The temporal models are typically categorized into three types: Recurrent Neural Networks (RNNs) [56], Temporal Convolution Networks (TCNs) [52, 57], and Transformers [58].
3 Method
3.1 GenSurg+
To enable effective video-language pre-training for robotic and laparoscopic surgery, we introduce GenSurg+, a large-scale dataset of surgical videos paired with descriptive captions. GenSurg+ builds upon the GenSurgery dataset [35], which was originally introduced as the largest publicly available dataset of general surgery videos.
3.2 Dataset Creation Pipeline
We began with the original GenSurgery dataset, which contains 3,100 videos spanning 28 different surgical procedures and totaling 680 hours of content. Our dataset creation pipeline involved several key steps to refine and augment this initial corpus:
Audio Filtering.
We first filtered out 1,300 videos that lacked audio content, as audio is crucial for generating meaningful textual descriptions.
Transcript Extraction.
For the remaining 1,800 videos with audio, we employed the Whisper model [59] to extract speech transcripts. This step was necessary as many of the videos, due to their age, lacked reliable YouTube automatic captions.
Video Segmentation and Filtering.
We segmented the videos into 45-second clips, resulting in approximately 18,000 individual segments. To ensure the quality and relevance of our dataset, we further filtered these clips based on linguistic criteria. Specifically, we removed about 1,000 clips that contained either too few unique words or highly repetitive content. This step helped eliminate silent segments and portions with non-informative audio (e.g., background music or noise).
Caption Generation.
For the remaining 17,000 high-quality video clips, we generated descriptive captions using the GPT-4 language model [60]. We crafted a specialized prompt to ensure the captions were concise, informative, and tailored to the surgical domain. Please see appendix for the prompt. The complete pipeline for creating GenSurg+ is illustrated in Figure 2.
Figure 2: Overview of the GenSurg+ dataset creation pipeline.
Table 1: Comparison between GenSurg+ and SVL-Pretrain datasets.
GenSurg+ SVL-Pretrain [33]
Publicly Available $\checkmark$ $\times$
##\## Videos 1.3k 1.3k
##\## Clip-Caption Pairs 17k -
Total Duration 213 hours -
3.3 Dataset Statistics and Characteristics
The resulting GenSurg+ dataset comprises 17,000 45-second video clips, totaling 213 hours of high-quality surgical content paired with descriptive captions. As shown in Table 1, this makes GenSurg+ the largest publicly available dataset specifically designed for surgical video-language pre-training, offering a significant resource for research in this area.
GenSurg+ represents a significant step forward in enabling large-scale video-language pre-training for robotic and laparoscopic surgery. By bridging the gap between visual content and descriptive text in the surgical domain, this dataset lays the foundation for more advanced and generalizable AI models in surgical assistance and analysis.
Figure 3: Overview of the VidLPRO model architecture and configuration. The model employs a Vision Transformer (ViT) as the video encoder and BERT as the text encoder. The multimodal fusion module integrates visual and textual representations, while pre-training objectives such as Video-Text Contrastive Learning (VTC), Video-Text Matching (VTM), and Masked Language Modeling (MLM) ensure comprehensive learning of multimodal representations.
3.4 VidLPRO
The VidLPRO framework is based on the best practices outlined in a comprehensive framework for video-language pre-training, adapted to the specific needs of surgical video analysis.
3.4.1 Model Architecture
Our VidLPRO model consists of three main components: a Video Encoder (VE), a Text Encoder (TE), and a Multimodal Fusion Module (MFM). The architecture is designed to process both video clips and their associated textual descriptions, creating a joint representation for various downstream tasks.
Video Encoder (VE).
We employ a standard Vision Transformer, specifically ViT-B/16 [61], as our video encoder. The ViT model is enhanced with a divided space-time temporal attention mechanism inspired by TimeSformer [62] to effectively capture the temporal dynamics of surgical videos. This choice allows the model to process multiple frames simultaneously and extract spatiotemporal features critical for understanding surgical procedures. Given a video clip $C=\{f_{1},f_{2},...,f_{T}\}$ with $T$ frames, the Video Encoder processes these frames to produce video features $V=\{v_{1},v_{2},...,v_{T}\}$:
$u_{t}=P(f_{t})$ (1)
$V=\text{VE}(\{u_{t}+p^{v}_{t}\}_{t=1}^{T})$ (2)
Here, $P(\cdot)$ is a linear projection, and $p^{v}_{t}$ are learnable positional embeddings that encode both spatial and temporal information. The ViT encoder is initialized using BEiT [63] weights.
Text Encoder (TE).
For the text encoder, we utilize BERT [64], a robust and widely-used transformer model for natural language processing. BERT is responsible for encoding the textual descriptions accompanying the surgical videos, such as transcripts and captions. For a given text description $D=\{w_{1},w_{2},...,w_{L}\}$ with $L$ tokens, the Text Encoder produces word embeddings $W=\{e_{1},e_{2},...,e_{L}\}$:
$W=\text{TE}(D)$ (3)
The BERT encoder is initialized with BERTbase [64] weights.
Multimodal Fusion Module (MFM).
The Multimodal Fusion Module integrates the visual and textual representations. We adopt the video-to-text (V2T) multimodal fusion scheme, which uses cross-attention to inject video cues into the textual features. The MFM takes the video features $V$ and word embeddings $W$ as input and performs cross-modal fusion to produce joint video-language representations $H$:
$H=\text{MFM}([V+p^{v},[CLS],W+p^{w}])$ (4)
where $p^{v}$ and $p^{w}$ are positional embeddings for video and text respectively, and [CLS] is a special token for classification tasks. The output $H$ can be divided into $H=[H^{v},h^{c},H^{w}]$, corresponding to video, global, and text representations. Following previous works [65, 66, 67], we reuse the text encoder and integrate a cross-attention operation into each of the last few layers of the text encoder, positioning it between the Self-Attention and MLP.
3.4.2 Pretraining Objectives
We employ three pretraining objectives to learn robust multimodal representations:
Video-Text Contrastive Learning (VTC).
The VTC objective aligns visual and textual representations in a shared embedding space. For a batch of $N$ video-text pairs, we compute:
$\mathcal{L}_{\text{VTC}}=(\mathcal{L}_{\text{v2t}}+\mathcal{L}_{\text{t2v}})/2$ (5)
where
$\mathcal{L}_{\text{v2t}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(sim(g^{v}_{i},g^{w}_{i})/\tau)}{\sum_{j=1}^{N}\exp(sim(g^{v}_{i},g^{w}_{j})/\tau)}$ (6)
$\mathcal{L}_{\text{t2v}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(sim(g^{w}_{i},g^{v}_{i})/\tau)}{\sum_{j=1}^{N}\exp(sim(g^{w}_{i},g^{v}_{j})/\tau)}$ (7)
Here, $g^{v}$ and $g^{w}$ are global video and text features obtained by applying a projection layer to the [CLS] token representation, $sim(\cdot,\cdot)$ is cosine similarity, and $\tau$ is a temperature parameter.
Video-Text Matching (VTM).
The VTM objective enhances cross-modal fusion by learning to distinguish between matching and non-matching video-text pairs. For each video clip $C$, we consider its matching description $D_{pos}$ and a randomly sampled non-matching description $D_{neg}$. We compute:
$s_{pos}=Q(h^{c}_{pos}),\quad s_{neg}=Q(h^{c}_{neg})$ (8)
$\mathcal{L}_{\text{VTM}}=-\mathbb{E}[\log(\sigma(s_{pos}))+\log(1-\sigma(s_{neg}))]$ (9)
where $Q(\cdot)$ is a linear layer, $h^{c}$ is the [CLS] token representation, and $\sigma(\cdot)$ is the sigmoid function.
Masked Language Modeling (MLM).
The MLM objective enhances the model’s understanding of surgical terminology. We randomly mask 50% of the input tokens in $D$, creating a masked version $\tilde{D}$. The model then predicts the original tokens:
$w^{\prime}_{i}=R(h^{w}_{\tilde{w}_{i}})$ (10)
$\mathcal{L}_{\text{MLM}}=-\mathbb{E}\left[\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\log P(w_{i}|w^{\prime}_{i})\right]$ (11)
where $R(\cdot)$ is a linear layer, $\mathcal{M}$ is the set of masked token indices, and $P(w_{i}|w^{\prime}_{i})$ is the probability of the correct token given the model’s prediction.
The full pre-training objective of VidLPRO is:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{VTC}}+\lambda_{2}\mathcal{L}_{\text{VTM}}+\lambda_{3}\mathcal{L}_{\text{MLM}}$ (12)
4 Experiments
To evaluate the effectiveness of our VidLPRO framework, we conducted extensive experiments on zero-shot surgical phase recognition tasks. We chose this task as it represents a challenging and clinically relevant application of video-language models in the surgical domain. Our experiments were designed to assess the generalizability and robustness of the representations learned by VidLPRO across different datasets and surgical procedures.
4.1 Pretraining Setup
We pre-trained VidLPRO on the GenSurg+ dataset. For each 45-second clip, we sampled 4 frames to capture temporal information while maintaining computational efficiency. Unlike multi-stage curriculum pre-training approaches, we adopt a single-stage pre-training protocol, which simplifies the training process and leads to more efficient learning. The pretraining was conducted using 4 NVIDIA A100 GPUs, allowing for efficient processing of the large-scale dataset. The video and text encoders were initialized with BEiT [63] and BERTbase [64] weights, respectively. More implementation details can be found in Table 2.
Table 2: Pre-training settings.
Configuration Value
Optimizer AdamW
Optimizer betas {0.9, 0.95}
Base learning rate 1e-4
Weight decay 0.02
Learning rate schedule Cosine schedule
Warm-up epochs 1
Epochs 50
Batch Size 256
Temperature $\tau$ 0.07
Loss weights $\lambda_{1}=\lambda_{2}=\lambda_{3}$ = 1
Table 3: Zero-shot Surgical Phase Recognition Performance On Cholec80 and AutoLaparo Datasets.
Model Cholec80 [14] AutoLaparo [36]
Accuracy (%) F1 Score (%) Accuracy (%) F1 Score (%)
MIL-NCE [68] 7.8 7.3 9.9 7.9
CLIP [31] 30.8 13.1 17.4 9.1
SurgVLP [33] 34.7 24.4 21.3 16.6
HecVL [34] 41.7 26.3 23.3 18.9
VidLPRO 57.1 (+15.4) 32.1 (+5.8) 42.5 (+19.2) 31.4 (+12.5)
4.2 Zero-Shot Surgical Phase Recognition
To evaluate the zero-shot capabilities of VidLPRO, we focused on two widely used datasets for surgical phase recognition: Cholec80 [14] and AutoLaparo [36]. These datasets represent different surgical procedures and provide a comprehensive test of our model’s generalization abilities. Cholec80 [14] consists of 80 videos of cholecystectomy procedures annotated with 7 surgical phases. AutoLaparo [36] contains 21 videos of laparoscopic hysterectomy procedures, divided into 7 phases.
To ensure a fair comparison with previous work, we adapted the class prompts used in SurgVLP [33] and HecVL [34] to better align with our caption-based pretraining approach. We used GPT-4 to transform the transcript-like class prompts into caption-like prompts, using the same prompt template employed for generating our pretraining captions. This process ensures that the evaluation prompts match the style and content of our pretraining data while maintaining the essential information about each surgical phase. The caption-like class prompts used for Cholec80 and AutoLapro datasets can be found in the appendix.
We split the videos from both datasets into 45-second clips, ensuring that each clip contains a single surgical phase. We then sampled 4 frames per clip, mirroring our pretraining setup. For zero-shot classification, we used the pretrained text encoder to extract representations of the class prompts and the video encoder to obtain representations of the video clips. The classification was performed by measuring the cosine similarity between the class prompt representations and the video clip representations, assigning each clip to the class with the highest similarity score.
4.3 Results and Comparison
We compared the performance of VidLPRO against several baselines, including SurgVLP [33], HecVL [34], and general-domain models like CLIP [31] and MIL-NCE [68] reported in HecVL [34]. Table 3 summarizes the results on both Cholec80 and AutoLaparo datasets.
The results demonstrate that VidLPRO achieves state-of-the-art zero-shot performance on both datasets, significantly outperforming previous surgical VLP methods. Notably, the general-domain models CLIP and MIL-NCE, which were pretrained on conventional computer vision datasets, perform poorly on these surgical tasks. The strong zero-shot performance of VidLPRO across two different datasets and different surgical procedures such as cholecystectomy and hysterectomy demonstrates the generalizability of the video-language representations learned by our model. These underscores the importance of domain-specific surgical pretraining and highlights the potential of VidLPRO as a foundation model for surgical video understanding.
Table 4: Ablation study results on the effect of the number of frames sampled per clip for zero-shot surgical phase recognition on Cholec80 and AutoLaparo datasets.
Frames Sampled Per Clip Cholec80 [14] AutoLaparo [36]
Accuracy (%) F1 Score (%) Accuracy (%) F1 Score (%)
1 50.9 28.8 41.2 30.7
4 57.1 32.1 42.5 31.4
8 57.9 32.2 43.8 33.0
16 59.1 33.7 43.3 32.4
32 60.1 33.8 44.0 33.4
45 61.0 (+19.3) 33.8 (+7.5) 44.8 (+21.5) 34.6 (+15.7)
HecVL [34] 41.7 26.3 23.3 18.9
4.4 Ablation Study on Number of Frames
To further understand the impact of design choices in VidLPRO, we conducted ablation studies focusing on the number of frames per clip used during inference. These experiments aim to identify the optimal configuration for balancing performance and computational efficiency during the zero-shot surgical phase recognition task. We evaluated VidLPRO’s performance using 1, 4, 8, 16, 32, and 45 frames per clip during inference. This range allows us to understand how the model’s performance scales with increased temporal information. Table 4 presents the results of these experiments on both Cholec80 and AutoLaparo datasets.
Experiments show that increasing the number of frames during inference generally leads to improved performance. This is expected, as more frames provide a richer representation of the surgical procedure, allowing for more accurate phase recognition. As we increase the number of sampled frames, the performance continues to improve. The margin of improvement becomes even larger when sampling 45 frames, showing that VidLPRO can effectively leverage additional temporal context when available. However, it’s important to note that the performance gains come with increased computational cost.
Given these trade-offs, we recommend using 4 frames for inference as a balanced configuration for most applications. With 4 frames, VidLPRO still significantly outperforms previous state-of-the-art methods while maintaining reasonable computational requirements. Notably, VidLPRO achieves state-of-the-art performance even when using only a single frame during inference, highlighting the robustness of the pre-trained representations.
5 Conclusion
This paper proposes VidLPRO, a novel video-language pre-training framework for surgical videos which first align the unimodal video and language representations before fusing them using multimodal module. Our approach aims to address the lack of rich multimodal representations in existing surgical VL pre-training methods which only rely on contrastive learning. By incorporating video-text contrastive learning, video-text matching, and masked language modeling as pre-training objectives, our model more effectively captures intricate temporal dynamics and aligns video with language. Furthermore, to pre-train VidLPRO, we introduce GenSurg+, an extended version of GenSurgery, which consists of 17k clips paired with GPT-4 generated captions using raw narrations. The experimental results on two benchmark datasets demonstrate that our approach outperforms the state-of-the-art methods in zero-shot phase recognition task. Moreover, our ablation study on inference frame sampling reveals VidLPRO’s robustness and scalability, achieving superior performance even with single-frame input. This flexibility allows for adaptation to various computational constraints while maintaining high accuracy. Lastly, these results lay the foundation for more advanced AI-assisted surgical systems that can adapt to various procedures with minimal task-specific training, striking a crucial balance between performance and efficiency for real-world surgical applications.
6 Limitations and Broader Impacts
Limitations.
Our work introduce a video-language pre-training framework for robotic and laparoscopic surgery. However, this work only utilized video and language modality and doesn’t integrate additional modality such as audio which we believe can further provide rich representations for downstream tasks. Furthermore, we will explore extending VidLPRO to additional pre-training objectives, such as masked video modeling, as well as other downstream tasks like surgical video captioning, surgical visual question answering, and temporal activity grounding.
Broader Impacts.
Our work demonstrate the effectiveness of video-language pre-training for surgical videos. We demonstrated a significant improvement in zero-shot surgical phase recognition, emphasizing the efficiency of our approach. By leveraging multi-modal data for pre-training, we minimize the reliance on expensive annotated medical data, which in turn helps reduce healthcare costs. Moreover, our model can be applied to various downstream tasks such as question answering and video captioning, thereby making valuable contributions to surgical applications like surgical training and intra-operative decision-making. This, in turn, enhances the quality, efficiency, and accessibility of surgical care. Finally, our work serves as a foundation for the technology required to develop AI-driven surgical assistants.
Appendices
7 Prompt for GenSurg+ caption
The prompt is designed to capture the essential surgical information while maintaining a professional and coherent tone.
"Generate a concise and informative caption that summarizes the main points of the narration. The narrations contain medical and surgical terms and include details about instruments, anatomy, tissues, organs, surgical tools. Make sure you don’t miss these in the generated captions. Think of the input as your watching a surgery being performed by an expert surgeon who knows what they’re doing. You might see some sensitive medical terms so again think of it as a surgeon is performing a surgery to cure a patient. Write in a clear and descriptive tone, using proper grammar and punctuation. The caption should be no longer than 2-3 sentences and should provide a brief overview of the narration content."
8 Class Prompts for phase labels
The caption-like textual prompts for Cholec-80 and AutoLapro are shown in Table 5 and 6 respectively.
Table 5: Caption-like textual prompts used for zero-shot surgical phase recognition task on the Cholec80 dataset.
Class Label Class Prompt
Preparation “Surgeon prepares for surgery by inserting trocars into the patient’s abdominal cavity.”
Calot Triangle Dissection “Surgeon employs grasper and hook during Calot triangle dissection, manipulating gallbladder to reveal hepatic triangle, cystic duct, and cystic artery."
Clipping Cutting "Surgeon utilizes clipper to secure cystic duct and artery, followed by precise dissection using scissors."
Gallbladder Dissection "Surgeon utilizes a hook to dissect the connective tissue during the dissection phase, separating the gallbladder from the liver."
Gallbladder Packaging "Surgeon secures the removed gallbladder in the specimen bag during the packaging phase of the procedure."
Cleaning Coagulation "Surgeon employs suction and irrigation techniques to maintain a clear surgical field during the clean and coagulation phase, simultaneously coagulating bleeding vessels."
Gallbladder Retraction "Surgeon expertly handles the specimen bag during the retraction phase, carefully extracting it from the trocar."
Table 6: Caption-like textual prompts used for zero-shot surgical phase recognition task on the AutoLaparo dataset.
Class Label Class Prompt
Preparation "During the preparation stage of a laparoscopic hysterectomy, the surgical team ensures all necessary instruments and equipment are sterilized and ready. The patient is anesthetized and positioned to optimize access to the pelvic area. Surgeons then make small incisions in the abdomen to insert the laparoscope and other surgical tools."
Dividing Ligament and Peritoneum "In this stage, surgeons carefully divide the ligament and peritoneum to access and isolate the uterus. This involves using specialized surgical instruments to delicately separate these tissues while preventing damage to surrounding organs. The division is done to create a clear surgical field and to facilitate the safe removal of the uterus."
Dividing Uterine Vessels and Ligament "At this stage, surgeons focus on meticulously severing the uterine vessels and ligaments. This is critical to control blood flow and prepare the uterus for removal. Specialized surgical tools are employed to ensure precision and minimize the risk of bleeding, ensuring a clear view and safe access to the target structures."
Transecting the Vagina "During this critical stage, the vagina is carefully transected at its connection to the cervix, using precise surgical techniques to ensure clean and controlled cuts. This step is essential for the complete removal of the uterus and cervix. Surgeons take extra precautions to maintain the integrity of the vaginal wall and surrounding tissues."
Specimen Removal "In the specimen removal stage, the excised uterus and any associated tissues are carefully extracted through one of the abdominal incisions. Surgeons may use a specialized bag to contain the specimen and minimize the risk of contamination. This step marks the completion of the critical surgical removal process, transitioning to closure and recovery procedures."
Suturing "In the suturing stage, the surgical sites, including the vagina and abdominal incisions, are meticulously closed using sutures. This is done to promote proper healing and prevent infection. Surgeons use various suturing techniques to ensure that the closures are secure and that the tissue alignment promotes optimal healing."
Washing "During the washing stage, the surgical area is thoroughly irrigated with sterile solutions to cleanse any debris and reduce the risk of post-operative infection. This step ensures that all remaining tissues are flushed and clear, providing a clean environment for the healing process to begin."
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims presented in our abstract and introduction accurately reflect the contributions and scope of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discussed the limitations of our work in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We don’t have any theoretical result and the related assumptions and proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provided all the implementation details to reproduce our results in Section 4.1.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We have provided all the implementation details to reproduce our experiments in Section 4.1. We also used a dataset that is publicly accessible.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provided all the experimental details, including hyperparameters in Section 4.1. We also conducted ablation studies necessary to understand the results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We reported zero-shot performance in Table 5 and Table 6 which don’t have error bars in the final results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Yes, we provided the compute resources in the Section 4.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Please see section 6 on broader impacts of our paper.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: There are no such risks in our paper.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have properly cited all the relevant assets related to models and data.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We haven’t introduced any new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We have used publicly available dataset, so our paper does not involve human subjects or crowdsourcing experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We have used publicly available dataset, so our paper does not involve human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.